STRATEGIES FOR COMPLEX MIXTURE ANALYSIS IN BROADBAND MICROWAVE SPECTROSCOPY.

\underline{AMANDA L. STEBER}, JUSTIN L. NEILL, MATTHEW T. MUCKLE, AND BROOKS H. PATE, Department of Chemistry, University of Virginia, McCormick Rd., P.O. Box 400319, Charlottesville, VA 22904; D.F. PLUSQUELLIC, Optical Technology Division, NIST, Gaithersburg, MD 20899-8441; V. LATTANZI, S. SPEZZANO, AND M.C. MCCARTHY, Harvard-Smithsonian Center for Astrophysics, 60 Garden St., Cambridge, MA 02138, and School of Engineering & Applied Sciences, Harvard University, 29 Oxford St., Cambridge MA 02138.

Broadband microwave spectra often contain overlapping spectra from a large number of species in the sample mixture, whether in the study of conformational isomers, molecular complexes, reaction products from reactive molecular sources (e.g., electrical discharge), or analysis of chemical mixtures. In these experiments, the identification of individual spectra in the full spectrum through pattern recognition becomes difficult when there is a high density of transitions. Strategies for extracting individual spectra from broadband measurements are discussed. Two approaches for microwave-microwave double resonance spectroscopy have been evaluated. One uses a transition-by-transition screening in a narrowband cavity spectrometer to identify an unknown spectrum and has a time advantage from the increased sensitivity of cavity spectroscopy. The second double-resonance approach uses a broadband spectral editing approach that gives a multiplex advantage in the detection. Both of these experimental techniques are combined with computer-aided assignment algorithms to make the spectral assignment in a minimum of double-resonance observations. The performance of spectral analysis solely using computer-aided assignment is also evaluated. The potential for fully automated spectral decomposition of the broadband spectrum of a complex mixture will be described.